OpenStudy (anonymous):

Help with this calculus question please. what is lim x->1 (2^x)*tan((pi*x)/2)-2*tan(x*(pi*x)/2))=??

OpenStudy (anonymous):

Will you convert it into eq using the equation section below the reply??

OpenStudy (anonymous):

Can you draw that equation out..

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1^-} 2^{x} \tan( \frac{ \pi x}{ 2 })-2\tan(x (\frac{ \pi x}{ 2 })) =?\]

OpenStudy (anonymous):

try substituting x-1=h

OpenStudy (anonymous):

first plug in 1- for x. we get positive infinity{1} - positive infinity{2}. to find out which infinity is bigger, take the limit of the ratio of the two, as x->1 =2^(x-1)*tan(pi x/2)/tan(pi x^2/2) =tan(pi x/2)/tan(pi x^2/2) =sin(pi x/2)cos(pi x^2/2)/(cos(pi x/2)sin(pi x^2/2)) =cos(pi x^2/2)/cos(pi x/2). let y=x-1 =cos(pi (y^2+2y+1)/2)/cos(pi (y+1)/2) =sin(y^2/2+y)/sin(y/2) use the squeeze thm sin(y)=y as y->0 =(y^2/2+y) / (y/2) =2+y =2 2 positive infinity-positive infinity = positive infinity

OpenStudy (anonymous):

I have got the answer by using l hospital. its positive infinity right. but thanks guys for the help.